OpenStudy (anonymous):

Solve for x: −3|2x + 6| = −12

OpenStudy (jhannybean):

begin by dividing each side by -3 so you get |2x+6| on one side of the equation

OpenStudy (anonymous):

x = 1 and x = 5 x = −1 and x = −5 x = −9 and x = 3 No solution those are the answers and i got 1 and 3

OpenStudy (anonymous):

-1 and 3

OpenStudy (jhannybean):

|2x + 6| = 4 -(2x + 6) = 4 2x + 6 = 4 -2x - 6 = 4 2x = 4 - 6 -2x = 10 2x = -2 x = 10/-2 x = -2/2

OpenStudy (anonymous):

so it would be -1 or -5 right?

OpenStudy (anonymous):

Thank you!

OpenStudy (jhannybean):

no problem :)